# Topological Obstructions to Dynamical Convexity

### v1

@@ Abstract.
We study the topological obstructions of dynamical convexity on contact manifolds focusing on fillability by cotangent bundles and subcritical surgeries. Using links to algebraic geometry, we motivate and define a stronger version of dynamical convexity, and investigate the topology of these manifolds. More precisely, we show that strongly dynamically convex contact manifolds cannot arise as a unit cotangent bundle $(ST^{*}M,\lambda_{std})$ of a closed manifold $M$ and in particular that dynamically convex contact manifolds cannot arise as the unit cotangent bundle $(ST^{*}M,\lambda_{std})$ of a simply connected manifold $M$. We also show obstructions to dynamical convexity that arises from studying different kinds of subcritical surgeries.
@@ 1. Introduction
A key notion for the dynamics on a contact boundary, introduced by Hofer, Wysocky and Zehnder to generalize the concept of affine convexity is dynamical convexity. In particular, they proved that the boundary of any affine convex domain in $\mathbb{R}^{2n}$ is dynamically convex. It was proven recently that the converse is not true [chaidez20223d, dardennes2024symplectic, chaidez2025ruelle].
The assumption of being dynamically convex is key in many results in dynamics, for instance, [hofer1999characterization, gutt2016minimal, abreu2017multiplicity, ginzburg2021dynamical, hryniewicz2022global],and for existence of some invariants, for instance in [hutchings2017cylindrical, zhou2021symplectic, chaidez2024contact]. We also highlight its relations to algebraic geometry, motivated by [mclean2016reeb]. The question we study in this paper is
@@
Let $\phi^{\alpha}$ be the flow generated by the Reeb vector field. We denote the set of closed orbits of this flow, called Reeb orbits by $\mathcal{P}(\alpha)$. Given a Reeb orbit $\gamma\in\mathcal{P}(\alpha)$, we can associate a number to $\gamma$ based on how much the flow of $R_{\alpha}$ winds around $\gamma$. This number is called the Conley Zehnder index of $\gamma$, $\mu_{CZ}(\gamma)$. More precisely, in a symplectic trivialization of $\gamma^{\star}TM$, the flow $\phi^{\alpha}$ gives a path of symplectic matrices to which we can associate a number; see [conley1984morse, robbin1993maslov, gutt2012conley].
A priori, the Conley Zehnder index depends on the choice of trivialization. This is not the case if the Reeb orbit is contractible and $c_{1}(\xi)=0.$ We can also have a well-defined Conley Zehnder index (independent of the trivialization) valued in $\mathbb{Q}$ denoted by $\mu^{\mathbb{Q}}_{CZ}$ if $c_{1}(\xi)$ is torsion. This latter index admits a lower semi-continuous extension $\mu^{\mathbb{Q}}_{LCZ}$ of $\mu^{\mathbb{Q}}_{CZ}$ [mclean2016reeb].
@@
It was proved in [hofer1999characterization], that if a contact 3-manifold admits a dynamically convex contact form, then the contact structure is tight and the second fundamental group of the manifold vanishes, thereby constraining its topology. Recently, [kwon2024dynamically] proved that if a contact manifold is dynamically convex, and simply connected, then it is contactomorphic to the standard sphere, provided the manifold admits a flexible Weinstein filling. This gives a characterisation of standard contact spheres.
@@
We prove that a large class of examples cannot admit such contact forms. Most notably, unit cotangent bundles of simply connected manifolds cannot admit dynamically convex contact forms.
@@
The unit cotangent bundle $ST^{*}M$ of a simply connected closed manifold $M$ cannot be dynamically convex.
Dynamical convexity is a condition defined only on contractible Reeb orbits. In [mclean2016reeb], Mclean showed that the Reeb dynamics on a the link of a singularity of an affine variety is related to an algebro-geometric invariant of the singularity called the minimal discrepancy. Using Mclean’s results, a conjecture in birational geometry concerning this invariant naturally leads to a stronger version of dynamical convexity which we call strong dynamical convexity (Definition 4.12). Motivated by these results we want to also study the topology of strongly dynamically convex contact manifolds.
We recall some definitions and results from [mclean2016reeb]. Given an irreducible affine variety $V$ of complex dimension $n$ in $\mathbb{C}^{N}$ with an isolated singularity at 0 (including the case where $V$ is smooth), the intersection of $V$ with a sphere $S^{2N-1}(\epsilon)$ centered at 0 for small enough $\epsilon$, is a 2n12n-1 differentiable manifold $V\cap S^{2N-1}(\epsilon). LL$ naturally inherits a contact structure $\xi_{L}=TL\cap i(TL)$.
The singularity of $V$ at 0 is called smooth if $V$ is smooth at 0 and topologically smooth if the link $L$ is diffeomorphic to $S^{2n-1}$. Note that if a singularity is smooth then it is topologically smooth. Mumford [mumford1961topology] proved that in complex dimension $2$, a singularity is smooth if and only if it is topologically smooth. However, in complex dimension $3$ and higher there are examples of non-smooth singularities which are topologically smooth. Therefore, the smoothness of the singularity can be determined by the diffeomorphism type of the link $L$ only in dimension 2.
@@
(Mclean[mclean2016reeb]) Given $H^{1}(L,\mathbb{Q})=0$, and $V$, an affine normal variety with an isolated numerically $\mathbb{Q}$-Gorenstein singularity, then
@@
for all Reeb orbits $\gamma$. ($M,\xi$) is said to be strongly dynamically convex [^1] This definition of strong dynamical convexity is different from the one introduced in [ginzburg2021dynamical] if there exists a strongly dynamically convex contact form $\alpha$ with $\ker\alpha=\xi$.
@@
An immediate observation in the direction of 1.10 is that if a contact manifold admits a contact form $\alpha$, such that $inf_{\{\gamma\in\mathcal{P}(\alpha)\}}\mu^{\mathbb{Q}}_{LCZ}(\gamma)\geq n+2$, then it cannot admit fillings $W$ with vanishing symplectic homology, in particular flexible fillings. This follows from the long exact sequence in symplectic homology 3.5 and the fact that $SH^{+}_{n+1}(W)=0$ which would force $H^{0}(W)=0$, a contradiction.
@@
Conjecture 1.6 is proven for toric varieties and up to complex dim $3$. Mclean uses 1.5 and the proof of Shokurov’s conjecture in complex dim $3$ to show that a singularity is smooth if and only if it is contactomorphic to the standard sphere in complex dimension $3$. Therefore strongly dynamically convex contact manifolds that arise as links are not just diffeomorphic but contactomorphic to the standard sphere up to real dimension $5$. We can thus strengthen our Question 1.9 to ask whether strongly dynamically convex manifolds are contactomorphic to standard spheres? Note that this is related to the problem of charaterizing $(S^{2n-1},\xi_{std})$. Since strong dynamical convexity and dynamical convexity coincide on a sphere, this would imply that the assumption of flexible fillability can be removed from the assumptions of [kwon2024dynamically, theorem A].
Question 1.10 would imply that strong dynamical convexity is the “limit case” for highest minimal index. A proof of 1.9 for links of singularities would thus prove a strong implication of conjecture 1.6.
@@
A consequence of theorem 1.13 is that simply connected unit cotangent bundles cannot be dynamically convex. Note that this is not the same as 1.3, which states that unit cotangent bundles over simply connected manifolds cannot be dynamically convex.
Another application of Theorem 1.13 is to check conjecture 1.6 for varieties with links of singularities that are contactomorphic to unit-cotangent bundles. A toy example is given by the affine variety $V(z_{1}^{2}+z_{2}^{2}+\cdots+z_{n}^{2})$ with singularity at origin whose link is contactomorphic to $(ST^{*}S^{n-1},\xi_{std}).$ [kwon2016brieskorn].
In a positive step towards Question 1.9, we have
@@
If M is dynamically convex, then by [hofer1999characterization, Theorem 1.5] $\xi$ is tight. By Poincaré’s theorem $M$ is diffeomorphic to a sphere $S^{3}$. The uniqueness of tight contact structure on $S^{3}$ [eliashberg1992contact] implies $M$ is contactomorphic to the standard tight contact sphere $(S^{3},\xi_{std}).$
Corollary 1.16.
@@
Note that $H^{1}(M,\mathbb{Q})=0.$ The universal coefficients theorem implies that $H_{1}(M,\mathbb{Q})=0$ and Poincaré duality implies $H_{2}(M,\mathbb{Q})=0$. Therefore, we have that $M$ is a rational homology sphere. ∎
Note that there exists examples of dynamically convex contact manifolds which are not spheres. We provide some of these examples in the next section.
In another direction, we looked at building dynamically convex manifold using flexible/subcritical surgeries (see [weinstein1991contact, cieliebak2002handle, fauck2016handle, lazarev2020contact]). This did not present new examples but obstructed large classes of manifolds from being dynamically convex, and hence strongly dynamically convex .
@@
Throughout the paper, when we refer to surgery, we use the framework of contact surgery as in [weinstein1991contact, fauck2016handle, geiges2008introduction] with the same framings.
@@
In section 2, we recall the definitions of dynamical and strong dynamical convexity,providing examples and counterexamples of these manifolds. In section 3, we review symplectic homology needed for the rest of the paper. We then prove theorem 1.13 and 1.3 using Serre spectral sequences and Viterbo’s isomorphism in section 4. In section 5, we prove the list of obstructions to dynamical convexity using subcritical surgery, more specifically theorems 1.18,1.19,1.20,1.21,1.22.
@@ Acknowledgements
I am indebted to both of my advisors, Jean François Barraud and Jean Gutt for all the discussions and advice and for their meticulous help with the exposition of this article. I would also like to acknowledge the financial support provided by EUR-MINT at the Institut de Mathématiques de Toulouse.
@@ 2. Dynamical Convexity
Given a path in Sp(2n2)Sp(2n-2) that starts at $id$ and ends in a matrix $A\in Sp(2n-2)$ such that $det(A-id)\neq 0$, one can associate an integer called Conley Zehnder index $\mu_{CZ}.$(cf[mclean2016reeb],[gutt2012conley] for details). The Conley Zehnder index $\mu_{CZ}$ has a lower semicontinous extension in Sp(2n2)Sp(2n-2) called $\mu_{LCZ}$cf [mclean2016reeb].
@@
In order to have a well defined canonical grading, we can assume that $c_{1}(\xi)$ is torsion , which suffices to define rational Conley Zehnder index $\mu^{\mathbb{Q}}_{CZ}$ of contractible Reeb orbits taking values in $\mathbb{Q}$. When $H^{1}(M,\mathbb{Q})=0$, we can obtain a well defined $\mu^{\mathbb{Q}}_{CZ}$ for not just contractible orbits but all orbits. (see[mclean2016reeb, Lemma 4.34.3],[li2025kahler] for details) .
@@
We say that a Reeb orbit $\gamma$ of period $T$ of contact form $\alpha$ is non-degenerate if $det(d{\phi^{\alpha}}_{T}-id)\neq 0.$ We say that $\alpha$ is non-degenerate if all of its Reeb orbits are non-degenerate. In [mclean2016reeb], $\mu^{\mathbb{Q}}_{LCZ}$ of an orbit is defined to be the infimum of non-degenerate nearby pertubations. Therefore $\mu^{\mathbb{Q}}_{LCZ}(\gamma)$ coincides with $\mu^{\mathbb{Q}}_{CZ}(\gamma)$ when $\gamma$ is non-degenerate. Therefore we can equivalently define dynamical convexity by using $\mu^{\mathbb{Q}}_{CZ}$ instead of $\mu^{\mathbb{Q}}_{LCZ}$ by restricting to non-degenerate contact forms.
@@
The definition of dynamical convexity that is prevalent in literature, for instance in [kwon2024dynamically][hofer1999characterization], usually includes the condition that the first chern class $c_{1}(\xi)$=0 instead of $c_{1}(\xi,\mathbb{Q})=0$. Note however that by 2.1, we have that normal definition is the same as ours when $c_{1}(\xi)=0$.
@@
2.5 would imply that lens spaces $L(p,q)$ with the standard contact structures they inherit from $S^{3}$ are dynamically convex. However, note that lens spaces arise as links of singularities[michel2020topology], therefore conjecture 1.6 which has been proved for links in dimension 3 would imply that they are not strongly dynamically convex.
@@
Note that strongly dynamically convex 33- manifolds are necessarily rational homology spheres by 1.16, thereby apart from lens spaces, a natural candidate to examine is the Poincare homology sphere $(P,\xi)$ with its standard contact structure. However $P$ being a brieskorn manifold arising as the link of a Brieskorn polynomial implies that its not strongly dynamically convex by the proof of 1.6 upto complex dimension $2$.
Definition 2.8 ([bowden2022tight]).
@@
[bowden2022tight, Proposition $3.2$] states that every contact manifold ($M^{2n-1},\xi$) with rational first chern class and supporting a contact form with no contractible orbit of conley zehnder index 4n4-n is algebraically tight. Algebraic tightness of a contact manifold implies tightness [casals2019geometric] . ∎

### v2

@@ Abstract.
We study the topological obstructions of dynamical convexity on contact manifolds focusing on fillability by cotangent bundles and subcritical surgeries. Using links to algebraic geometry, we motivate and define a stronger version of dynamical convexity, and investigate the topology of these manifolds. More precisely, we show that strongly dynamically convex contact manifolds cannot arise as a unit cotangent bundle $(ST^{*}M,\lambda_{std})$ of a closed manifold $M$ and in particular that simply connected dynamically convex contact manifolds cannot be filled by cotangent bundles. We demonstrate that dynamical convexity can be used to recover homotopy groups of topologically simple fillings with vanishing symplectic homology. We also show obstructions to dynamical convexity that come from studying different kinds of subcritical surgeries.
@@ 1. Introduction
A key notion for the dynamics on a contact boundary, introduced by Hofer, Wysocki and Zehnder to generalize the concept of affine convexity is dynamical convexity. In particular, they proved that the boundary of any affine convex domain in $\mathbb{R}^{2n}$ is dynamically convex. It was proven recently that the converse is not true [CE22, DGZ24, CE25].
The assumption of being dynamically convex is key in many results in dynamics, for instance, [HWZ99, GK16, AM17, GM21, HSS22],and for existence of some invariants, for instance in [HN17, ZHO21, CHA24]. We also highlight its relations to algebraic geometry, motivated by [MCL16]. The question we study in this paper is
@@
Let $\phi^{\alpha}$ be the flow generated by the Reeb vector field. We denote the set of closed orbits of this flow, called Reeb orbits by $\mathcal{P}(\alpha)$. Given a Reeb orbit $\gamma\in\mathcal{P}(\alpha)$, we can associate a number to $\gamma$ based on how much the flow of $R_{\alpha}$ winds around $\gamma$. This number is called the Conley Zehnder index of $\gamma$, $\mu_{CZ}(\gamma)$. More precisely, in a symplectic trivialization of $\gamma^{\star}TM$, the flow $\phi^{\alpha}$ gives a path of symplectic matrices to which we can associate a number; see [CZ84, RS93, GUT12].
A priori, the Conley Zehnder index depends on the choice of trivialization. This is not the case if the Reeb orbit is contractible and $c_{1}(\xi)=0.$ We can also have a well-defined Conley Zehnder index (independent of the trivialization) valued in $\mathbb{Q}$ denoted by $\mu^{\mathbb{Q}}_{CZ}$ if $c_{1}(\xi)$ is torsion. This latter index admits a lower semi-continuous extension $\mu^{\mathbb{Q}}_{LCZ}$ of $\mu^{\mathbb{Q}}_{CZ}$ [MCL16].
@@
It was proved in [HWZ99], that if a contact 3-manifold admits a dynamically convex contact form, then the contact structure is tight and the second fundamental group of the manifold vanishes, thereby constraining its topology. Recently, [KO24] proved that if a contact manifold is dynamically convex, and simply connected, then it is contactomorphic to the standard sphere, provided the manifold admits a flexible Weinstein filling. This gives a characterisation of standard contact spheres.
@@
We prove that a large class of examples cannot admit such contact forms. Most notably, simply connected unit cotangent bundles of manifolds cannot admit dynamically convex contact forms.
@@
The unit cotangent bundle $ST^{*}M$ of a closed manifold $M$ cannot be dynamically convex if it is simply connected .
Dynamical convexity is a condition defined only on contractible Reeb orbits. In [MCL16], Mclean showed that the Reeb dynamics on a the link of a singularity of an affine variety is related to an algebro-geometric invariant of the singularity called the minimal discrepancy. Using Mclean’s results, a conjecture in birational geometry concerning this invariant naturally leads to a stronger version of dynamical convexity which we call strong dynamical convexity (Definition 1.7). Motivated by these results we want to also study the topology of strongly dynamically convex contact manifolds.
We recall some definitions and results from [MCL16]. Given an irreducible affine variety $V$ of complex dimension $n$ in $\mathbb{C}^{N}$ with an isolated singularity at 0 (including the case where $V$ is smooth), the intersection of $V$ with a sphere $S^{2N-1}(\epsilon)$ centered at 0 for small enough $\epsilon$, is a 2n12n-1 differentiable manifold $V\cap S^{2N-1}(\epsilon). LL$ naturally inherits a contact structure $\xi_{L}=TL\cap i(TL)$.
The singularity of $V$ at 0 is called smooth if $V$ is smooth at 0 and topologically smooth if the link $L$ is diffeomorphic to $S^{2n-1}$. Note that if a singularity is smooth then it is topologically smooth. Mumford [MUM61] proved that in complex dimension $2$, a singularity is smooth if and only if it is topologically smooth. However, in complex dimension $3$ and higher there are examples of non-smooth singularities which are topologically smooth. Therefore, the smoothness of the singularity can be determined by the diffeomorphism type of the link $L$ only in dimension 2.
@@
(Mclean[MCL16]) Given $H^{1}(L,\mathbb{Q})=0$, and $V$, an affine normal variety with an isolated numerically $\mathbb{Q}$-Gorenstein singularity, then
@@
for all Reeb orbits $\gamma$. ($M,\xi$) is said to be strongly dynamically convex if there exists a strongly dynamically convex contact form $\alpha$ with $\ker\alpha=\xi$.
Remark 1.8.
This definition of strong dynamical convexity is different from the one introduced in [GM21].
@@
An immediate observation in the direction of 1.11 is that if a contact manifold admits a contact form $\alpha$, such that $inf_{\{\gamma\in\mathcal{P}(\alpha)\}}\mu^{\mathbb{Q}}_{LCZ}(\gamma)\geq n+2$, then it cannot admit fillings $W$ with vanishing symplectic homology, in particular flexible fillings. This follows from the long exact sequence in symplectic homology 3.5 and the fact that $SH^{+}_{n+1}(W)=0$ which would force $H^{0}(W)=0$, a contradiction.
@@
Conjecture 1.6 is proven for toric varieties and up to complex dim $3$. Mclean uses 1.5 and the proof of Shokurov’s conjecture in complex dim $3$ to show that a singularity is smooth if and only if it is contactomorphic to the standard sphere in complex dimension $3$. Therefore strongly dynamically convex contact manifolds that arise as links are not just diffeomorphic but contactomorphic to the standard sphere up to real dimension $5$. We can thus strengthen our Question 1.10 to ask whether strongly dynamically convex manifolds are contactomorphic to standard spheres? Note that this is related to the problem of charaterizing $(S^{2n-1},\xi_{std})$. Since strong dynamical convexity and dynamical convexity coincide on a sphere, this would imply that the assumption of flexible fillability can be removed from the assumptions of [KO24, theorem A].
Question 1.11 would imply that strong dynamical convexity is the “limit case” for highest minimal index. A proof of 1.10 for links of singularities would thus prove a strong implication of conjecture 1.6.
@@
A consequence of theorem 1.14 is that simply connected unit cotangent bundles cannot be dynamically convex. Note that unit cotangent bundles of simply connected manifolds are simply connected if $n\geq 3$ by using the long exact sequence of homotopy groups of a fibration. For $n=2$, the unit cotangent bundle of $S^{2}$ which is $\mathbb{RP}^{3}$ is dynamically convex, however not strongly dynamically convex.
Another application of Theorem 1.14 is to check conjecture 1.6 for varieties with links of singularities that are contactomorphic to unit-cotangent bundles. A toy example is given by the affine variety $V(z_{1}^{2}+z_{2}^{2}+\cdots+z_{n}^{2})$ with singularity at origin whose link is contactomorphic to $(ST^{*}S^{n-1},\xi_{std}).$ [KV16]. Other examples include unit cotangent bundles of manifolds considered in [MCL18, Theorem $1.1$].
In a positive step towards Question 1.10, we have
@@
If M is dynamically convex, then by [HWZ99, Theorem 1.5] $\xi$ is tight. By Poincaré’s theorem $M$ is diffeomorphic to a sphere $S^{3}$. The uniqueness of tight contact structure on $S^{3}$ [ELI92] implies $M$ is contactomorphic to the standard tight contact sphere $(S^{3},\xi_{std}).$
Proposition 1.17.
@@
For a manifold to be strongly dynamically convex, we impose the condition $H^{1}(M,\mathbb{Q})=0$ in the definition for having a well defined Conley Zehnder index for non-contractible orbits. The universal coefficients theorem implies $H_{1}(M,\mathbb{Q})=0$ since $H^{1}(M,\mathbb{Q})=0$ and Poincaré duality implies $H_{2}(M,\mathbb{Q})=0$. Therefore, $M$ is a rational homology sphere. ∎
Note that there exists examples of dynamically convex contact manifolds which are not spheres but are quotients of spheres. We provide some of these examples in the next section.
[KO24] proved that dynamically convex contact manifolds with fillings having vanishing symplectic homology are homology spheres and when $\pi_{1}=0$, homeomorphic to spheres. We prove that we can obtain homotopical information of such fillings without the condition on $\pi_{1}$ by working with symplectic homology of covering spaces. More precisely, we prove the following statement.
Theorem 1.18 (Theorem 3.19).
Let $M$ be a dynamically convex contact manifold such that there exists a topologically simple filling $W^{\prime}$ of $M$ with $SH(W^{\prime})=0$. Then for any topologically simple filling $W^{\prime}$ of $M$, $\pi_{k}(W)\cong\pi_{k}(M)$ for 1<k<2n11<k<2n-1 if $\pi_{1}(\partial W)\rightarrow\pi_{1}(W)$ is surjective.
In another direction, we looked at building dynamically convex manifold using flexible/subcritical surgeries (see [WEI91, CIE02, FAU16, LAZ20]). This did not present new examples but obstructed large classes of manifolds from being dynamically convex, and hence strongly dynamically convex .
@@
Throughout the paper, when we refer to surgery, we use the framework of contact surgery as in [WEI91, FAU16, GEI08] with the same framings.
@@
In section 2, we recall the definitions of dynamical and strong dynamical convexity,providing examples and counterexamples of these manifolds. In section 3, we review symplectic homology and prove theorem 3.19 using symplectic homology with local coefficients. We then prove theorem 1.14 and 1.3 using Serre spectral sequences and Viterbo’s isomorphism in section 4. In section 5, we prove the list of obstructions to dynamical convexity using subcritical surgery, more specifically theorems 1.20,1.21,1.22,3.19,1.24.
@@ Acknowledgements
I am indebted to both of my advisors, Jean François Barraud and Jean Gutt for all the discussions and advice and for their meticulous help with the exposition of this article. I would like to thank Alberto Abbondandolo for valuable comments regarding the first version of the paper. I would also like to acknowledge the financial support provided by EUR-MINT at the Institut de Mathématiques de Toulouse.
@@ 2. Dynamical Convexity
Given a path in Sp(2n2)Sp(2n-2) that starts at $id$ and ends in a matrix $A\in Sp(2n-2)$ such that $det(A-id)\neq 0$, one can associate an integer called Conley Zehnder index $\mu_{CZ}.$(cf[MCL16],[GUT12] for details). The Conley Zehnder index $\mu_{CZ}$ has a lower semicontinous extension in Sp(2n2)Sp(2n-2) called $\mu_{LCZ}$ cf [MCL16].
@@
In order to have a well defined canonical grading, we can assume that $c_{1}(\xi)$ is torsion , which suffices to define rational Conley Zehnder index $\mu^{\mathbb{Q}}_{CZ}$ of contractible Reeb orbits taking values in $\mathbb{Q}$. When $H^{1}(M,\mathbb{Q})=0$, we can obtain a well defined $\mu^{\mathbb{Q}}_{CZ}$ for not just contractible orbits but all orbits as the set of homotopy classes of trivializations are determined by $H^{1}(M,\mathbb{Q})$ . (see[MCL16, Lemma 4.34.3],[LZ25] for details) .
@@
We say that a Reeb orbit $\gamma$ of period $T$ of a contact form $\alpha$ is non-degenerate if $det(d{\phi^{\alpha}}_{T}-id)\neq 0.$ We say that $\alpha$ is non-degenerate if all of its Reeb orbits are non-degenerate. In [MCL16], $\mu^{\mathbb{Q}}_{LCZ}$ of an orbit is defined to be the infimum of non-degenerate nearby pertubations. Therefore $\mu^{\mathbb{Q}}_{LCZ}(\gamma)$ coincides with $\mu^{\mathbb{Q}}_{CZ}(\gamma)$ when $\gamma$ is non-degenerate. Therefore we can equivalently define dynamical convexity by using $\mu^{\mathbb{Q}}_{CZ}$ instead of $\mu^{\mathbb{Q}}_{LCZ}$ by restricting to non-degenerate contact forms.
@@
The definition of dynamical convexity that is prevalent in literature, for instance in [KO24][HWZ99], usually includes the condition that the first chern class $c_{1}(\xi)$=0 instead of $c_{1}(\xi,\mathbb{Q})=0$. Note however that by 2.1, we have that normal definition is the same as ours when $c_{1}(\xi)=0$.
@@
2.5 would imply that lens spaces $L(p,q)$ with the standard contact structures they inherit from $S^{3}$ are dynamically convex. However, note that lens spaces arise as links of singularities[MIC20], therefore conjecture 1.6 which has been proved for links in dimension 3 implies that they are not strongly dynamically convex.
@@
Note that strongly dynamically convex 33- manifolds are necessarily rational homology spheres by 1.17, thereby apart from lens spaces, a natural candidate to examine is the Poincare homology sphere $(P,\xi)$ with its standard contact structure. However $P$ being a brieskorn manifold arising as the link of a Brieskorn polynomial implies that its not strongly dynamically convex by the proof of 1.6 upto complex dimension $2$.
Definition 2.8 ([BGM+22]).
@@
[BGM+22, Proposition $3.2$] states that every contact manifold ($M^{2n-1},\xi$) with rational first chern class and supporting a contact form with no contractible orbit of conley zehnder index 4n4-n is algebraically tight. Algebraic tightness of a contact manifold implies tightness [CMP19] . ∎